THE OPTICAL STARK SPECTRUM OF THE A^{3}Φ_{4} − X^{3}Φ_{4} BAND SYSTEM OF IRIDIUM MONOFLUO-RIDE, IrF

\underline{XIUJUAN ZHUANG} AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; COLAN LINTON, Center for Lasers, Atomic and Molecular Sciences and Physics Department, University of New Brunswick, Fredericton, NB Canada E3B 5A3.

Recently the New Brunswick group^{a} reported on the field-free detection and analysis of the A^{3}Φ_{4} − X^{3}Φ_{4} band system of IrF. Here we report on the analysis Q(4)(15922 cm^{−1}) branch feature of the (1,0) band of the ^{191}IrF isotopologue of that system recorded at field strengths of up to 3000 V/cm. The spectra are surprisingly complex at the achieved resolution of 40 MHz due to the presence of both the ^{191}Ir(I=3/2) and ^{19}F(I=1/2) magnetic hyperfine splitting. The determined permanent electric dipole moment, μ_{el}, for the X^{3}Φ_{4} state is compared with that recently determined ^{b} for the X^{3}Φ_{4} state of isovalent CoF. The trend in μ_{el} amongst the ground states of IrF, IrC and IrN ^{c}will be discussed. Finally, a simple molecular orbital correlation diagram will be used to rationalize the change in μ_{el} upon excitation from the X^{3}Φ_{4} to A^{3}Φ_{4} state.

aA.G. Adam; A.D. Granger; L.E. Downie; D.W. Tokaryk and C. Linton Can.J. Phys. 87 557, 2009.

bH. Wang; X. Zhaung; and T.C. Steimle J. Chem. Phys. 131 114315, 2009.

cA.J.Marr; M.E. Flores; and T.C. Steimle J. Chem. Phys. 104 8183, 1996.